# OmniEdit: Building Image Editing Generalist Models Through Specialist Supervision

### v1

@@ 1 Introduction
We curate an image editing benchmark Omni-Edit-Bench, which contains diverse images of different resolutions and diverse prompts that cover all the listed editing skills. We perform comprehensive automatic and human evaluation to show the significant boost of Omni-Edit over the existing baseline models like CosXL-Edit  (Boesel & Rombach, 2024), UltraEdit  (Zhao et al., 2024), etc. On automatic metrics like VIEScore (Ku et al., 2024), we can outperform all the existing approaches by a reasonable margin in terms of both perceptual quality and semantic consistency. We also performed a human evaluation and observed an overall improvement 20% over the best baseline editing model CosXL-Edit (Boesel & Rombach, 2024).
@@ 3 Learning with Specialist Supervision
Figure 2: Overview of the Omni-Edit training pipeline.
@@ 3.2 Constructing Specialist Models
Object Addition. We treat object addition as the inverse task of object removal. Specifically, for each pair of editing examples generated by the object removal specialist, we reverse the roles of the source and target images to create a new pair.
@@
Background Swap. We trained an image inpainting model to serve as the specialist $q_{\text{obj\_background\_swap}}$. We use a similar procedure as in the object replacement but use an inverse mask of the object to indicate the background and guide the inpainting.
Environment Modification. For environment modification, we use P2P pipeline to generate original and edited image. To ensure structural consistency between two images, we apply a mask of the foreground to maintain details in the foreground while changing the background.
@@ 3.3 Importance weighting
While the GPT-4o is an effective choice for this task, scoring large-scale datasets with millions of examples is extremely costly and time-consuming. Therefore, we employ knowledge distillation from GPT-4o to a smaller 8B model, InternVL2 (Chen et al., 2024b). For each task, we sample 50K data points and instruct GPT-4o to output both a score and a score rationale. We fine-tune InternVL2 on these GPT-4o-generated examples. After fine-tuning, InternVL2 performs as an ideal scoring function due to its smaller size and efficiency. A comparison of the model’s performance before and after fine-tuning is presented in the Appendix. Finally, we apply the fine-tuned InternVL2 model to filter data across a dataset with millions of samples. Only examples with a score of $\geq 9$ are retained, resulting in a curated training dataset of 775K examples. We visualize InternVL2’s response as a scoring function before and after fine-tuning in Figure 3. We observe that fine-tuning InternVL2 on GPT-4o’s response effectively turns InternVL2 into a realiable scoring function and it can identify unusual distortions or unsuccessful edit that does not follow the editing instruction. Additional dataset statistics are detailed in the Appendix.
@@ 5 Experiments
Omni-Edit Training Dataset. We constructed the training dataset $\mathcal{D}$ by sampling high-resolution images with a minimum resolution of 1 megapixel from the LAION-5B (Schuhmann et al., 2022) and OpenImageV6 (Kuznetsova et al., 2020) databases. The images cover a range of aspect ratios including 1:1, 2:3, 3:2, 3:4, 4:3, 9:16, and 16:9. For the task of object swap, we employed a specialist model to generate 1.5 million entries. We then applied InternVL2 for importance weighting, retaining samples with scores of 9 or higher, resulting in a dataset of 150K entries for this task. Similarly, we generate 250k-1M samples for each task, then keep the top 10% as the final dataset. The final training dataset comprises 775K entries, with detailed information provided in Appendix 6.
@@
Omni-Edit implementation details. The Omni-Edit model is built upon Stable diffusion 3 Medium(Esser et al., 2024) with EditNet architecture. The stable diffusion 3 has 24 DiT layers. Each layer has a corresponding EditNet layer. We train Omni-Edit on the 775K Omni-Edit training dataset for 2 epochs on a single node with 8 H100 GPUs.
@@ 5.2 Ablation Study
Ablation study on the importance sampling. We study a baseline that utilizes the same architecture as Omni-Edit, but instead of applying importance scoring and filtering, we sample 775K examples directly from the 5M pre-filtering dataset as specified in Table 6 and compare it with Omni-Edit. As shown in Table 4, we observe a significant decrease in VIEScores for both PQ and SC metrics.
@@
Table 4: Ablation on importance sampling.
Models VIEScore (GPT4o) VIEScore (Gemini)
Omni-Edit 8.38 6.66 6.98 7.06 5.82 5.78
Omni-Edit w/o importance sampling 6.20 2.95 3.30 6.40 1.80 2.25
Table 5: Ablation on Omni-Edit architecture design.
Models VIEScore (GPT4o) VIEScore (Gemini)
Omni-Edit 8.38 6.66 6.98 7.06 5.82 5.78
Omni-Edit- ControlNet - TextControl 6.45 4.70 4.89 6.50 4.35 4.48
Omni-Edit- ControlNet 6.35 4.60 4.75 6.40 4.25 4.35
@@ 6 Related Work
Image Editing via Generation. Editing real images according to specific user requirements has been a longstanding research challenge (Crowson et al., 2022; Liu et al., 2020; Zhang et al., 2023; Shi et al., 2022; Ling et al., 2021). Since the introduction of large-scale diffusion models, such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2024), significant progress has been made in tackling image editing tasks. SDEdit (Meng et al., 2022) introduced an approach that adds noise to the input image at an intermediate diffusion step, followed by denoising guided by the target text description to generate the edited image. Subsequent methods, such as Prompt-to-Prompt (Hertz et al., 2023) and Null-Text Inversion (Mokady et al., 2023), have focused on manipulating attention maps during intermediate diffusion steps for image editing. Other techniques like Blended Diffusion (Avrahami et al., 2022) and DiffEdit (Couairon et al., 2022) utilize masks to blend regions of the original image into the edited output. More recently, the field has seen a shift towards supervised methods, such as InstructP2P (Brooks et al., 2023), HIVE (Zhang et al., 2024b), and MagicBrush (Zhang et al., 2024a), which incorporate user-written instructions in an end-to-end framework. Our work follows this direction to develop end-to-end editing models without inversion.
@@ Appendix A Appendix
Object Swap 1,500,000 150,000
Object Removal 1,000,000 100,000
Object Addition 1,000,000 100,000
Background Swap 500,000 50,000
Environment Change 500,000 100,000
Style Transfer 250,000 25,000
@@ A.1.8 Importance Sampling
We apply the importance sampling as described in Section 3.3. Example prompts that are provided to LMMs are shown in Figure 8 and 9. We compute the Overall score following (Ku et al., 2024) as the importance weight. After importance sampling, we obtain our training dataset described in Table 6.

### v2

@@ 1 Introduction
We curate an image editing benchmark Omni-Edit-Bench, which contains diverse images of different resolutions and diverse prompts that cover all the listed editing skills. We perform comprehensive automatic and human evaluation to show the significant boost of Omni-Edit over the existing baseline models like CosXL-Edit (Boesel & Rombach, 2024), UltraEdit (Zhao et al., 2024), etc.
@@ 2.2 Instruction-Based Image Editing in Supervised Learning
Figure 2: Overview of the Omni-Edit training pipeline. The pipeline consists of four stages: (1) task-specific specialist models are trained for diverse editing tasks; (2) these specialist models are used to generate a large, high-resolution, multi-aspect-ratio dataset; (3) a cost-efficient distilled large multi-modal model (LMM) assigns importance weights to each pair of image-editing data; and (4) the final generalist model is trained on the weighted dataset.
@@ 3.2 Constructing Specialist Models
Object Addition. We treat object addition as the inverse task of object removal.
@@
Background Swap. We use a similar procedure as in the object replacement but use an inverse mask of the object to indicate the background and guide the inpainting.
Environment Modification. For environment modification, we use P2P pipeline to generate original and edited image.
@@ 3.3 Importance weighting
While the GPT-4o is an effective choice for this task, scoring large-scale datasets with millions of examples is extremely costly and time-consuming. Therefore, we employ knowledge distillation from GPT-4o to a smaller 8B model, InternVL2 (Chen et al., 2024b). For each task, we sample 50K data points and instruct GPT-4o to output both a score and a score rationale. We fine-tune InternVL2 on these GPT-4o-generated examples. After fine-tuning, InternVL2 performs as an ideal scoring function due to its smaller size and efficiency. A comparison of the model’s performance before and after fine-tuning is presented in the Appendix. Finally, we apply the fine-tuned InternVL2 model to filter data across a dataset with millions of samples. Only examples with a score of $\geq 9$ are retained, resulting in a curated training dataset of 1.2M examples. We visualize InternVL2’s response as a scoring function before and after fine-tuning in Figure 3. We observe that fine-tuning InternVL2 on GPT-4o’s response effectively turns InternVL2 into a realiable scoring function and it can identify unusual distortions or unsuccessful edit that does not follow the editing instruction. Additional dataset statistics are detailed in the Appendix.
@@ 5 Experiments
Omni-Edit Training Dataset. We constructed the training dataset $\mathcal{D}$ by sampling high-resolution images with a minimum resolution of 1 megapixel from the LAION-5B (Schuhmann et al., 2022) and OpenImageV6 (Kuznetsova et al., 2020) databases. The images cover a range of aspect ratios including 1:1, 2:3, 3:2, 3:4, 4:3, 9:16, and 16:9. For the task of object swap, we employed a specialist model to generate 1.5 million entries. We then applied InternVL2 for importance weighting, retaining samples with scores of 9 or higher, resulting in a dataset of 150K entries for this task. Similarly, we generate 250k-1M samples for each task, then keep the top 10% as the final dataset. The final training dataset comprises 1.2M entries, with detailed information provided in Appendix 4.
@@
Omni-Edit implementation details. The Omni-Edit model is built upon Stable diffusion 3 Medium(Esser et al., 2024) with EditNet architecture. The stable diffusion 3 has 24 DiT layers. Each layer has a corresponding EditNet layer. We train Omni-Edit on the 1.2M Omni-Edit training dataset for 2 epochs on a single node with 8 H100 GPUs.
@@ 5.2 Ablation Study
Ablation study on the importance sampling. We study a baseline that utilizes the same architecture as Omni-Edit, but instead of applying importance scoring and filtering, we sample 1.2M examples directly from the 5M pre-filtering dataset as specified in Table 4 and compare it with Omni-Edit. As shown in Table 6, we observe a significant decrease in VIEScores for both PQ and SC metrics.
@@ 6 Related Work
Image Editing via Generation. Editing real images according to specific user requirements has been a longstanding research challenge (Crowson et al., 2022; Liu et al., 2020; Zhang et al., 2023; Shi et al., 2022; Ling et al., 2021; Wasserman et al., 2024; Ju et al., 2024). Since the introduction of large-scale diffusion models, such as Stable Diffusion (Rombach et al., 2022; Podell et al., 2024), significant progress has been made in tackling image editing tasks. SDEdit (Meng et al., 2022) introduced an approach that adds noise to the input image at an intermediate diffusion step, followed by denoising guided by the target text description to generate the edited image. Subsequent methods, such as Prompt-to-Prompt (Hertz et al., 2023) and Null-Text Inversion (Mokady et al., 2023), have focused on manipulating attention maps during intermediate diffusion steps for image editing. Other techniques like Blended Diffusion (Avrahami et al., 2022) and DiffEdit (Couairon et al., 2022) utilize masks to blend regions of the original image into the edited output. More recently, the field has seen a shift towards supervised methods, such as InstructP2P (Brooks et al., 2023), HIVE (Zhang et al., 2024b), and MagicBrush (Zhang et al., 2024a), which incorporate user-written instructions in an end-to-end framework. Our work follows this direction to develop end-to-end editing models without inversion.
@@ Appendix A Appendix
Object Swap 1,500,000 180,000
Object Removal 1,000,000 200,000
Object Addition 1,000,000 200,000
Background Swap 500,000 180,000
Environment Change 500,000 160,000
Style Transfer 250,000 50,000
@@ A.1.8 Importance Sampling
We apply the importance sampling as described in Section 3.3. Example prompts that are provided to LMMs are shown in Figure 6 and 7. We compute the Overall score following (Ku et al., 2024) as the importance weight. After importance sampling, we obtain our training dataset described in Table 4.
@@ A.2 Additional Evaluation Result
Table 6: Ablation on importance sampling.
Models VIEScore (GPT4o) VIEScore (Gemini)
Omni-Edit 8.38 6.66 6.98 7.06 5.82 5.78
Omni-Edit w/o importance sampling 6.20 2.95 3.30 6.40 1.80 2.25
Table 7: Ablation on Omni-Edit architecture design.
Models VIEScore (GPT4o) VIEScore (Gemini)
Omni-Edit 8.38 6.66 6.98 7.06 5.82 5.78
Omni-Edit- ControlNet - TextControl 6.45 4.70 4.89 6.50 4.35 4.48
Omni-Edit- ControlNet 6.35 4.60 4.75 6.40 4.25 4.35